what is the original amount of Po if 25.0 grams after 5 half lives?
1
Expert's answer
2019-01-10T03:47:15-0500

The radiactive half-life gives a pattern of reduction to half in any successive half-life period.

A=A2-t/T

T-is half-life period

As t=5T then

A=A2-5T/T

25=A2-5

25=A/32

A= 32*25

A=800

Answer: 800 g